OpenStudy (anonymous):

could you solve this for me please? y=(x^2-5x+8)^6. can i expand through with the exponent 6?? how do i solve this.

ganeshie8 (ganeshie8):

you wanto differentiate it ?

ganeshie8 (ganeshie8):

if so, use chain rule

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

show me please

Parth (parthkohli):

Do you know the Chain Rule?

OpenStudy (anonymous):

yeah...but dis looks soo confusing

Parth (parthkohli):

OK. Let's see: The first function we see is x^2 - 5x + 8. The second function that is acting on the first function is (...)^6.

OpenStudy (anonymous):

ok

Parth (parthkohli):

Recall the formula for the Chain Rule.\[\dfrac{d}{dx}f(g(x)) = f'(g(x)) \times g'(x)\]

Parth (parthkohli):

So what is \(f'(g(x))\)?

Parth (parthkohli):

Hmm, do you have any thoughts so far? Are you able to follow?

Parth (parthkohli):

\[f(g(x))\]The \(g\) is the "inside" function and the \(f\) is another function which is "outside" the inside function.

OpenStudy (anonymous):

ok

Parth (parthkohli):

So here, the \(g\) is \(x^2 - 5x + 8\) because it's the "inside" function. The second function \(f\) is the "something to the power six" function because it's acting on the inside function.

OpenStudy (anonymous):

f′(g(x))=6(2x-5)^5

Parth (parthkohli):

Hmm, you leave the \(x^2 - 5x + 8\) in that as it is. You're only differentiating the outside function.

OpenStudy (anonymous):

okay

Parth (parthkohli):

\[f'(g(x)) = 6(x^2 - 5x + 8)^5 \]\[g'(x) = 2x - 5\]

OpenStudy (anonymous):

so d/dxf(g(x))=f′(g(x))×g′(x) ==6(x^2-5x+8)^5 * (2x-5)

OpenStudy (anonymous):

thanks a loads